OpenStudy (anonymous):

The length of the rectangle is 3 less than twice it's width. Find the dimensions given the perimeter is 84 ft.

OpenStudy (anonymous):

Perimeter = 2* length + 2 * width --- (P = 2L + 2W) P = 84 L = 2W - 3 now we sub P = 2L + 2W 84 = 2(2W - 3) + 2W 84 = 4W - 6 + 2W 84 = 6W - 6 84 + 6 = 6W 90 = 6W 90/6 = W 15 = W L = 2W - 3 L = 2(15) - 3 L = 30 - 3 L = 27 check.. P = 2L + 2W 84 = 2(27) + 2(15) 84 = 54 + 30 84 = 84 (correct) So the width (W) = 15 and the length (L) = 27

OpenStudy (anonymous):

Thanks!! :D